OpenStudy (dls):

A man of height 1.5m walks away along a straight line from a lamp post which is 6m high at the rate of 3m/s.Then find (i)Rate of increase of length of its shadow. (ii)Rate of change of end of the shadow.

OpenStudy (dls):

 1.5/x=6/y x=3y.. dx/dt=3y.dy/dt dx/dt is given 3 so dy/dt=1 maybe? not sure about 2nd part though